OpenStudy (anonymous):

Given a regular decagon, find the measures of the angles formed by (a) two consecutive radii and (b)a radius and a side of the polygon.

OpenStudy (anonymous):

a) reqd angle is 360/10=36

OpenStudy (anonymous):

b) the reqd angle is (180-36)/2=144/2=72